OpenStudy (anonymous):

find the area of rhombus with p=100 and a diagonal of 48

OpenStudy (anonymous):

what is p?

OpenStudy (anonymous):

Area of a rhombus is given by: \(\frac 12 \times\) Product of the diagonals.

OpenStudy (anonymous):

perimeter

OpenStudy (anonymous):

what do i do with the perimeter

OpenStudy (anonymous):

If P is the perimeter then P/4 is the length of the one side right?

OpenStudy (anonymous):

yeah 25

OpenStudy (anonymous):

i don't understand what to do with that

OpenStudy (anonymous):

Okay, There is a property relating side and the two diagonals of a rhombus: side\( ^2 \) = ( \(\frac12\times \) one diagonal)\(^2\) + (\(\frac12\times \) other diagonal)\(^2\) So, you can use this to find the length of the other diagonal and then use it to find the area.

OpenStudy (anonymous):

so one diagonal would equal 96

OpenStudy (anonymous):

One diagonal is 48 which is given and the other should be 14. Area = 336 sq units.

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

how did you get 14?

OpenStudy (anonymous):

Set x = the other diagonal, now using the formula